A $5,000 bond with a coupon rate of 5.7​% paid semiannually has four years to maturity and a yield to maturity of 6.8​%.

If interest rates fall and the yield to maturity decreases by​ 0.8%, what will happen to the price of the​ bond?
Answer #1

Calculating Price of bond when Yield to maturity is 6.8%

Face value = $5000,

As coupons are paid semi annually, therefore

Semi annual coupon payment = (Coupon rate x Face value) / 2 = (5.7% x 5000) / 2 = 285 / 2 =$142.50

No of half years to maturity = 2 x no of years to maturity = 2 x 4 = 8

Semi annual yield to maturity = Yield to maturity / 2 = 6.8% / 2 = 3.4%

Price of a bond can be found out finding present value of semi annual coupon payments and face value payment by discounting at semi annual yield to maturity. We can find the present value or price using PV function in excel

Formula to be used in excel: =PV(rate,nper,-pmt,-fv)

н 5000! 142.50| 3.40%! 96 Calculating price of bond 97 Face value (fv) 98 Semi Annual Coupon payment (pmt) 99 Semi annual Yie

Using PV function in excel, we get Price of bond when Yield to maturity is 6.8% = $4810.1747

Calculating Price of bond when Yield to maturity decreases by 0.8%

New Yield to maturity = 6.8% - 0.8% = 6.0%

New Semi annual Yield to maturity = New Yield to maturity / 2 = 6% / 2 = 3%

Now we can use PV function in excel to calculate new price of bond

Formula to be used in excel: =PV(rate,nper,-pmt,-fv)

H 96 CalculatingNew price of bond 97 Face value (fv) 98 Semi Annual Coupon payment (pmt) 99 New Semi annual Yield to maturity

Using PV function in excel, we get new price of bond = $4947.3523

We can see that due decrease in Yield to maturity or interest rates, price of bond rises

Increase in Price of bond = New price - Old price = 4947.3523 - 4810.1747 = 137.1776 = $137.18

Answer: Price of bond will rise by $137.18